# No-Rank Tensor Decomposition Using Metric Learning

### v1

@@ 3.2. Triplet Loss with Regularization
To prevent dimensional collapse and encourage the model to use all available dimensions efficiently, we add a diversity penalty on the embedding correlation matrix:
$\mathcal{L}_{\text{div}}=\frac{1}{d(d-1)}\sum_{i\neq j}|\mathbf{C}_{ij}|,$ (6)
where $\mathbf{C}=\mathbf{Z}^{\top}\mathbf{Z}$ is the correlation matrix of the embeddings $\mathbf{Z}=[\mathbf{z}_{1},\ldots,\mathbf{z}_{n}]^{\top}$.
Following [wang2020understanding], we also promote a uniform distribution of embeddings on the unit sphere to avoid hubness and further improve generalization:
@@ 3.3. Locality Preservation Framework
To ensure that local neighborhoods in the original high-dimensional space are preserved in the embedding space, we introduce locality preservation objectives. Let $\mathcal{N}_{k}(\mathbf{x}_{i})$ denote the set of $k$-nearest neighbors of $\mathbf{x}_{i}$ in the original space. The goal is to ensure:
$\mathbf{x}_{j}\in\mathcal{N}_{k}(\mathbf{x}_{i})\Rightarrow f(\mathbf{x}_{j})\in\mathcal{N}_{k}(f(\mathbf{x}_{i})).$ (8)
@@
where $\mathcal{N}_{k}^{O}$ and $\mathcal{N}_{k}^{E}$ are neighborhoods in the original and embedding spaces, respectively, and $r_{O}(i,j)$ is the rank of $j$ in $\mathbf{x}_{i}$’s original neighborhood. Trustworthiness measures the prevalence of ”intruders” (points in the embedding neighborhood that were not in the original neighborhood):
@@
where $\mathcal{U}_{k}(\mathbf{x}_{i})$ is the set of intruders for point $i$.
@@
Unlike low-rank tensor decomposition methods (e.g., CP or Tucker), which seek a rank-$R$ approximation by minimizing reconstruction error $\|\mathcal{X}-\hat{\mathcal{X}}\|_{F}^{2}$, the approach learns embeddings without explicit rank constraints by directly optimizing for semantic similarity. Furthermore, while traditional tensor decompositions are limited to linear transformations, the use of a neural network encoder allows for the capture of complex non-linear patterns.
@@ 3.4. Network Architecture
The embedding function $f(\cdot)$ is implemented as a deep neural network. The encoder, which maps an input $\mathbf{x}$ to the unit sphere, has the following structure:
@@
The final $\ell_{2}$-normalization projects the embedding onto the unit sphere, which is compatible with the uniformity loss. A non-linear projection head can be added to enhance representation quality during training:
$\mathbf{p}=\mathbf{W}_{2}\cdot\text{ReLU}(\mathbf{W}_{1}\mathbf{z}).$ (17)
@@ 3.5. Convergence and Geometric Guarantees
We now establish theoretical guarantees for the proposed framework, focusing on optimization convergence and the geometric properties of the learned embedding space.
If the mapping $f$ is $L$-Lipschitz continuous, i.e.,
$\|f(\mathbf{x})-f(\mathbf{x}^{\prime})\|_{2}\leq L\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2},$
then local neighborhood relationships are preserved up to a factor of $L$, providing a theoretical guarantee for locality preservation.
For a metric learning model with Rademacher complexity $\mathfrak{R}_{n}$ [shalev2014understanding, mohri2018foundations], the generalization error $\mathcal{E}_{\text{gen}}$ is bounded with probability at least $1-\delta$ by:
@@
The gradients of the triplet loss with respect to the embeddings create a force field that pulls positive pairs together and pushes negative pairs apart [schroff2015facenet]. For an active triplet (where the loss is positive), the gradients are:
@@
almost surely, where $\mathcal{F}_{t}$ denotes the history up to time $t$. When minimized using stochastic gradient descent with a learning rate schedule $\{\eta_{t}\}$ satisfying the Robbins–Monro conditions $\sum_{t=1}^{\infty}\eta_{t}=\infty$ and $\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty$ [robbins1951stochastic], the parameter sequence $\{\theta_{t}\}$ satisfies
@@
Under the assumptions of Lemmas 1 and 2, the proposed metric learning framework:
@@
The theorem follows from combining the guarantees of the two lemmas. From Lemma 1, the optimization converges to a critical point $\theta^{*}$ where $\liminf_{t\to\infty}\mathbb{E}[\|\nabla_{\theta}\mathcal{L}_{\text{total}}(\theta_{t})\|^{2}]=0$. Under the strict saddle point assumption [^1] The strict saddle property requires that every saddle point has at least one direction of negative curvature. Under this condition, gradient-based methods with random initialization almost surely avoid saddle points [lee2016gradient]., gradient-based methods converge almost surely to local minima rather than saddle points. Thus, $\theta^{*}$ is a local minimum of $\mathcal{L}_{\text{total}}$ with high probability.
From Lemma 2, when the embedding function $f_{\theta^{*}}$ (parameterized by the locally optimal $\theta^{*}$) operates on separable data with margin $\gamma$, the resulting embedding space exhibits the semantic structure described in Lemma 2. Specifically (i) intra-class distances are bounded by $L\cdot\text{diam}(\mathcal{M}_{c})$, (ii) inter-class distances are lower-bounded by $\sqrt{\gamma}$.
@@
    The Lipschitz constant $L$ (from Lemma 2) controls how well local neighborhoods are preserved,
@@
In practice, the regularization terms $\mathcal{L}_{\text{div}}$ and $\mathcal{L}_{\text{uniform}}$ promote well-spread embeddings (supporting large $\gamma$), while $\mathcal{L}_{\text{local}}$ and $\mathcal{L}_{\text{global}}$ enforce neighborhood preservation (related to the Lipschitz property).
@@ 4.1. Evaluation Metrics
A higher ratio indicates that the embedding space successfully pulls same-class samples together and pushes different-class samples apart.
@@ 4.2. Visualization of High-Dimensional Embeddings
We employ three standard dimensionality reduction methods Principal Component Analysis (PCA) [doi:10.1080/14786440109462720] as a linear baseline that projects data onto the directions of maximal variance. t-Distributed Stochastic Neighbor Embedding (t-SNE) [maaten2008visualizing] as a non-linear technique that preserves local similarities by minimizing the Kullback-Leibler divergence [kullback1951information] between probability distributions in the high- and low-dimensional spaces:
@@
And Uniform Manifold Approximation and Projection (UMAP) [mcinnes2018umap] as a manifold learning technique that assumes the data is uniformly distributed on a Riemannian manifold. It constructs a topological representation and optimizes a low-dimensional equivalent.
@@ 4.3. Triplet Mining Strategy Analysis
The strategy for selecting triplets from the training data is critical for efficient and stable model convergence. We analyze two common strategies [schroff2015facenet, wang2019multi, hermans2017defense].
@@ 5.1. Metric Learning Performance in Face Recognition
We used two contrasting datasets (Table 1): the Labeled Faces in the Wild (LFW) [huang2008labeled] dataset, a medium-sized, imbalanced dataset representing a real-world challenge; and the Olivetti Faces [pedregosa2011scikit] dataset, a smaller, balanced dataset captured in a controlled environment.
@@
Figure 1. t-SNE visualization of face embeddings. Left to right: Olivetti (original data), Olivetti (metric learning), LFW (original data), LFW (metric learning). Metric learning produces distinct, identity-based clusters.
The practical impact of this clustering is evident in nearest-neighbor retrieval. In Figures 2 and 3, the nearest neighbors of a query image in the metric learning space are consistently of the same person, despite variations in pose and lighting. The small distances between anchor and positive pairs confirm the model successfully compacts same-identity samples.
Figure 2. Nearest-neighbor retrieval on the LFW dataset using the metric learning embedding. The model correctly identifies same-identity images as the closest neighbors, with small corresponding Euclidean distances.
Figure 3. Nearest-neighbor retrieval on the Olivetti Faces dataset using the metric learning embedding. The results demonstrate robust, identity-based retrieval.
The rank sensitivity analysis reveals a fundamental limitation of fixed-rank tensor decompositions for semantic clustering tasks. On the challenging LFW dataset, both CP and Tucker decompositions fail to achieve meaningful clustering across all ranks, with Silhouette scores remaining near zero or negative. This indicates that the low-rank constraints destroy the semantic structure necessary for identity separation. While Tucker decomposition shows a slight improvement with higher ranks (from -0.0689 at R5 to -0.0037 at R20), the gains are minimal. The Olivetti dataset, with its controlled conditions, shows more rank sensitivity, particularly for Tucker decomposition where performance improves substantially with higher ranks (0.1303 at R10 to 0.1845 at R20). However, even the best fixed-rank result (Tucker-R20: 0.1845) is significantly outperformed by the metric learning approach (0.8566), demonstrating that rank constraints inherently limit the ability to capture semantically meaningful representations, regardless of rank selection.

### v2

@@ 3.4. Triplet Loss with Regularization
The fundamental triplet loss was popularized in FaceNet for deep metric learning [40], and subsequent work has explored its optimization and sampling strategies [55]. To regularize embedding spaces, recent research has highlighted the importance of alignment and uniformity properties for robust representations [52], and proposed variants that additionally encourage diversity or decorrelation among features [33].
@@
To prevent dimensional collapse and encourage the model to use all available dimensions efficiently, we employ the diversity penalty defined in Proposition 3 where $\mathbf{C}=\mathbf{Z}^{\top}\mathbf{Z}$ as defined above.
Following [52], we also promote a uniform distribution of embeddings on the unit sphere to avoid hubness and further improve generalization:
@@ 3.5. Locality Preservation Framework
Unlike low-rank tensor decomposition methods (e.g., CP, Tucker, or t-SVD), which enforce global rank constraints by minimizing reconstruction error, the proposed framework prioritizes local semantic consistency. This shift enables flexible, non-linear embeddings that adapt to intrinsic data geometry rather than enforcing a fixed algebraic structure. To ensure that local neighborhoods in the original high-dimensional space are preserved in the embedding space, we introduce locality preservation objectives. Let $\mathcal{N}_{k}(\mathbf{x}_{i})$ denote the set of $k$-nearest neighbors of $\mathbf{x}_{i}$ in the original space. The objective is to encourage preservation of local neighborhoods, i.e.,
$\mathbf{x}_{j}\in\mathcal{N}_{k}(\mathbf{x}_{i})\;\Rightarrow\;f(\mathbf{x}_{j})\text{ remains close to }f(\mathbf{x}_{i}).$ (8)
@@
where $\mathcal{N}_{k}^{O}$ and $\mathcal{N}_{k}^{E}$ are neighborhoods in the original and embedding spaces, respectively, and $r_{O}(i,j)$ is the rank of $j$ in $\mathbf{x}_{i}$’s original neighborhood. Here, continuity is computed using a rank-weighted penalty consistent with neighborhood overlap measures commonly used in manifold learning. Trustworthiness measures the prevalence of ”intruders” (points in the embedding neighborhood that were not in the original neighborhood):
@@
where $\mathcal{U}_{k}(\mathbf{x}_{i})$ is the set of intruders for point $i$. These quantities are used as evaluation metrics and are not directly optimized due to their non-differentiability.
@@ 3.6. Network Architecture
The embedding function $f(\cdot)$ is realized as a deep neural network that maps inputs $\mathbf{x}$ onto the unit sphere. The encoder consists of $L$ fully connected layers with ReLU activations:
@@
The final $\ell_{2}$-normalization ensures that embeddings lie on the unit sphere, facilitating uniformity regularization. Optionally, a non-linear projection head
$\mathbf{p}=\mathbf{W}_{2}\,\text{ReLU}(\mathbf{W}_{1}\mathbf{z}),$ (17)
can be added to improve representation quality during training, similar to contrastive learning frameworks. By the universal approximation theorem [21], fully connected networks with sufficient width and depth can approximate any continuous mapping on compact domains, ensuring the encoder has enough capacity to capture complex non-linear relationships.
3.7. Model Complexity
The proposed framework has a space complexity determined by the encoder network. For an $L$-layer fully connected network with layer widths $\{h_{1},\dots,h_{L}\}$, input dimension $D$, and embedding dimension $d$, the total number of trainable parameters is $\mathcal{O}(Dh_{1}+\sum_{l=1}^{L-1}h_{l}h_{l+1}+h_{L}d)$. Unlike CP or Tucker decompositions, the method does not require storing tensor factors or specifying a rank. The time complexity for a single sample forward pass is of the same order, while training with triplets scales linearly with the number of triplets in a batch. Additional regularization terms contribute modest overhead: the diversity loss costs $\mathcal{O}(nd^{2})$ per batch, and locality-preserving losses cost $\mathcal{O}(nkd)$ where we assume $k$-nearest neighbors are precomputed in the input space and fixed during training. In contrast, traditional tensor decompositions scale with both tensor size and rank (e.g., $\mathcal{O}(R\prod_{m}I_{m})$ for CP), making the proposed approach more scalable for high-dimensional data while maintaining a flexible, rank-free embedding framework.
A schematic of the proposed no-rank metric learning framework for tensor data is dipicted in Figure 1 (a detailed illustration can be found in Appendix 7, Figure 12.)
[…]
Figure 1. Overview of the proposed no-rank metric learning framework. Tensor slices are embedded via a neural encoder and optimized with triplet-based metric learning, combined with geometric and diversity regularization, producing a semantically structured embedding space. The embedding space shows tight intra-class clustering and large inter-class separation. For a more detailed illustration of the method see Figure 12.
@@ 3.8. Convergence and Geometric Guarantees
To provide theoretical context, we leverage generalization bounds based on Rademacher complexity [42, 34]. Metric learning theory supports locality preservation under Lipschitz mappings [54]. Optimization guarantees for SGD under smooth objectives have been established in [8, 31]. Manifold geometry and neighborhood structure preservation have been studied in classical manifold learning literature [17, 44]. We now establish theoretical guarantees for the proposed framework, focusing on optimization convergence and the geometric properties of the learned embedding space.
If the mapping $f$ is $L$-Lipschitz, local neighborhoods are distorted by at most a factor proportional to $L$, providing a sufficient condition for approximate locality preservation.
For a metric learning model with Rademacher complexity $\mathfrak{R}_{n}$ [42, 34], the generalization error $\mathcal{E}_{\text{gen}}$ is bounded with probability at least $1-\delta$ by:
@@
The gradients of the triplet loss induce attractive forces between positive pairs and repulsive forces between negative pairs [40]. For an active triplet (where the loss is positive), the gradients are:
@@
almost surely, where $\mathcal{F}_{t}$ denotes the history up to time $t$. When minimized using stochastic gradient descent with a learning rate schedule $\{\eta_{t}\}$ satisfying the Robbins–Monro conditions $\sum_{t=1}^{\infty}\eta_{t}=\infty$ and $\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty$ [38], the parameter sequence $\{\theta_{t}\}$ satisfies
@@
Lemma 9 (Metric Preservation via Triplet Loss and Lipschitz Mapping).
Assume the data lies on a compact manifold $\mathcal{M}$ and is separable with margin $\gamma>0$, i.e., for all valid triplets $(a,p,n)$:
$\|\mathbf{z}_{a}-\mathbf{z}_{p}\|_{2}^{2}+\gamma\leq\|\mathbf{z}_{a}-\mathbf{z}_{n}\|_{2}^{2}.$
Let the embedding function $f:\mathcal{M}\to\mathbb{R}^{d}$ be $L$-Lipschitz:
$\|f(\mathbf{x})-f(\mathbf{y})\|_{2}\leq L\|\mathbf{x}-\mathbf{y}\|_{2}\quad\forall\mathbf{x},\mathbf{y}\in\mathcal{M}.$
Then, for any two points $\mathbf{x}_{i},\mathbf{x}_{j}\in\mathcal{M}$:
  1. (i)
    Intra-class preservation: if $\mathbf{x}_{i},\mathbf{x}_{j}$ belong to the same class component $\mathcal{M}_{c}$,
    $\|f(\mathbf{x}_{i})-f(\mathbf{x}_{j})\|_{2}\leq L\,d_{\mathcal{M}}(\mathbf{x}_{i},\mathbf{x}_{j})\leq L\,\text{diam}(\mathcal{M}_{c}),$
    so local semantic distances are preserved up to a factor $L$.
  2. (ii)
    Inter-class separation: if $\mathbf{x}_{i},\mathbf{x}_{j}$ belong to different classes,
    $\|f(\mathbf{x}_{i})-f(\mathbf{x}_{j})\|_{2}\geq\sqrt{\gamma},$
    providing a lower bound on distances across classes.
Consequently, the embedding $f$ defines an approximate isometry on the manifold with bounded distortion:
$\frac{1}{L}\|f(\mathbf{x}_{i})-f(\mathbf{x}_{j})\|_{2}\leq d_{\mathcal{M}}(\mathbf{x}_{i},\mathbf{x}_{j})\leq L\|f(\mathbf{x}_{i})-f(\mathbf{x}_{j})\|_{2}\quad\text{(for intra-class pairs)}.$
Proof.
Intra-class: For $\mathbf{x}_{i},\mathbf{x}_{j}\in\mathcal{M}_{c}$, Lipschitz continuity gives
$\|f(\mathbf{x}_{i})-f(\mathbf{x}_{j})\|_{2}\leq L\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}\leq L\,d_{\mathcal{M}}(\mathbf{x}_{i},\mathbf{x}_{j})\leq L\,\text{diam}(\mathcal{M}_{c}),$
preserving local distances up to factor $L$.
Inter-class: For $\mathbf{x}_{i},\mathbf{x}_{j}$ from different classes, consider a triplet with anchor $a=\mathbf{x}_{i}$, positive $p$ in the same class, and negative $n=\mathbf{x}_{j}$. The triplet loss implies
$\|\mathbf{z}_{a}-\mathbf{z}_{p}\|_{2}^{2}+\gamma\leq\|\mathbf{z}_{a}-\mathbf{z}_{n}\|_{2}^{2}\implies\|f(\mathbf{x}_{i})-f(\mathbf{x}_{j})\|_{2}\geq\sqrt{\gamma}.$
Combining the two results shows that $f$ approximately preserves intra-class manifold distances while enforcing inter-class separation, yielding a bounded-distortion embedding suitable for similarity-based tasks. ∎
@@
Under the assumptions of Lemmas 7, 8, and 9, the proposed metric learning framework:
@@
  3. (3)
    Preserves pairwise distances on the data manifold up to a bounded distortion factor, i.e., for any $\mathbf{x}_{i},\mathbf{x}_{j}\in\mathcal{M}$:
    $\frac{1}{L}\|\mathbf{z}_{i}-\mathbf{z}_{j}\|_{2}\leq d_{\mathcal{M}}(\mathbf{x}_{i},\mathbf{x}_{j})\leq L\|\mathbf{z}_{i}-\mathbf{z}_{j}\|_{2},$ (22)
    where $\mathbf{z}_{i}=f(\mathbf{x}_{i})$ and $d_{\mathcal{M}}$ is the geodesic distance on the manifold.
@@
The theorem follows from combining the guarantees of the two lemmas. From Lemma 7, the optimization converges to a critical point $\theta^{*}$ where $\liminf_{t\to\infty}\mathbb{E}[\|\nabla_{\theta}\mathcal{L}_{\text{total}}(\theta_{t})\|^{2}]=0$. Under the strict saddle point assumption [^1] The strict saddle property requires that every saddle point has at least one direction of negative curvature. Under this condition, gradient-based methods with random initialization almost surely avoid saddle points [31]., gradient-based methods converge almost surely to local minima rather than saddle points. Thus, $\theta^{*}$ is a local minimum of $\mathcal{L}_{\text{total}}$ with high probability.
From Lemma 8, when the embedding function $f_{\theta^{*}}$ (parameterized by the locally optimal $\theta^{*}$) operates on separable data with margin $\gamma$, the resulting embedding space exhibits the semantic structure described in Lemma 8. Specifically (i) intra-class distances are bounded by $L\cdot\text{diam}(\mathcal{M}_{c})$, (ii) inter-class distances are lower-bounded by $\sqrt{\gamma}$. Finally, due to Lemma 9, because $f$ is $L$-Lipschitz, the embedding approximately preserves pairwise distances on the manifold:
$\frac{1}{L}d_{\mathcal{M}}(\mathbf{x}_{i},\mathbf{x}_{j})\leq\|f(\mathbf{x}_{i})-f(\mathbf{x}_{j})\|_{2}\leq L\,d_{\mathcal{M}}(\mathbf{x}_{i},\mathbf{x}_{j}),$
providing a bounded distortion guarantee.
@@
    The Lipschitz constant $L$ (from Lemma 8 and the metric preservation lemma) controls how well local neighborhoods are preserved and bounds the distortion of pairwise distances on the data manifold,
@@
In practice, the regularization terms $\mathcal{L}_{\text{div}}$ and $\mathcal{L}_{\text{uniform}}$ promote well-spread embeddings (supporting large $\gamma$), while $\mathcal{L}_{\text{local}}$ and $\mathcal{L}_{\text{global}}$ enforce neighborhood preservation and approximate isometry (related to the Lipschitz property).
@@ 4.1. Evaluation Metrics
Since the Separation Ratio is scale-invariant, its absolute magnitude depends on the relative contraction of intra-class distances and expansion of inter-class distances rather than the embedding norm; therefore, comparisons are meaningful across methods trained under identical preprocessing and normalization. A higher ratio indicates that the embedding space successfully pulls same-class samples together and pushes different-class samples apart.
@@ 4.2. Visualization of High-Dimensional Embeddings
We employ three standard dimensionality reduction methods Principal Component Analysis (PCA) [14] as a linear baseline that projects data onto the directions of maximal variance. t-Distributed Stochastic Neighbor Embedding (t-SNE) [46] as a non-linear technique that preserves local similarities by minimizing the Kullback-Leibler divergence [30] between probability distributions in the high- and low-dimensional spaces:
@@
And Uniform Manifold Approximation and Projection (UMAP) [32] as a manifold learning technique that assumes the data is uniformly distributed on a Riemannian manifold. It constructs a topological representation and optimizes a low-dimensional equivalent.
@@ 4.3. Triplet Mining Strategy Analysis
The strategy for selecting triplets from the training data is critical for efficient and stable model convergence. We analyze two common strategies [40, 53, 18].
@@ 5. Results
To comprehensively evaluate the proposed no-rank metric learning framework, we compare its performance against representative methods from three major categories: (i) classical linear and manifold-based embeddings, (ii) tensor decomposition–based approaches with explicit rank selection, and (iii) deep learning–based representation and clustering models.
(i) Linear and Manifold-Based Baselines:
We include widely used dimensionality reduction techniques followed by k-means clustering, namely PCA, t-SNE, and UMAP. These methods serve as standard baselines for assessing clustering quality and neighborhood preservation in low-dimensional embeddings, but do not explicitly model tensor structure or semantic similarity.
(ii) Tensor Decomposition Methods:
To benchmark against classical tensor-based representations, we consider CP, Tucker, and t-SVD decompositions with varying ranks ($R=5,10,20$). For each method, the input tensor is decomposed into low-rank factors, vectorized representations are extracted, and k-means clustering is applied. These approaches require explicit rank selection and primarily optimize reconstruction fidelity rather than discriminative or semantic objectives.
(iii) Deep Learning Baselines:
We further compare against representative reconstruction-based deep clustering models, including Variational Autoencoders (VAE) [25] and Deep Embedded Clustering (DEC) [56]. These methods learn latent representations by optimizing reconstruction-driven objectives and perform clustering in the learned latent space, without explicitly enforcing metric or neighborhood constraints.
Transformer-Based Models:
Transformer architectures are evaluated separately in Subsection 5.5, as they introduce fundamentally different inductive biases, optimization dynamics, and computational costs compared to both tensor decompositions and autoencoder-based models. A dedicated analysis allows for a more focused comparison with the proposed metric learning framework.
In contrast to all baselines, the proposed method directly learns an embedding space optimized for semantic similarity using a triplet-based metric learning objective augmented with diversity, uniformity, and locality-preserving regularization. Unlike tensor decompositions, it does not impose explicit rank constraints, and unlike reconstruction-based deep models, it does not aim to reconstruct the original tensor. Instead, the embedding dimensionality acts as an implicit, optimization-driven notion of rank.
All methods are evaluated using identical preprocessing, consistent train–test splits, and matched embedding dimensionality where applicable. Clustering is performed using k-means for all baselines to ensure comparability. Performance is assessed using clustering compactness (Silhouette, Davies–Bouldin), separation (Separation Ratio), neighborhood preservation (Continuity, Trustworthiness), and clustering agreement metrics (ARI, NMI). Dataset-specific results are reported in the following subsections.
@@ 5.1. Metric Learning Performance in Face Recognition
We used two contrasting datasets (Table 1): the Labeled Faces in the Wild (LFW) [22] dataset, a medium-sized, imbalanced dataset representing a real-world challenge; and the Olivetti Faces [37] dataset, a smaller, balanced dataset captured in a controlled environment.
@@
The clustering performance of the proposed metric learning approach is a direct consequence of how the embedding is learned. Unlike classical dimensionality reduction and tensor decomposition methods, which aim to preserve variance or minimize reconstruction error, the proposed metric learning model explicitly optimizes the embedding geometry for semantic separability through the metric learning objective.
Specifically, the triplet loss enforces that samples of the same identity are embedded close together while pushing different identities far apart, producing a representation that is intrinsically cluster-friendly. In contrast, methods such as PCA, CP, Tucker, and t-SVD focus on low-rank approximation of the input data, and nonlinear techniques such as t-SNE and UMAP emphasize neighborhood preservation rather than global class separation. Consequently, these baselines yield embeddings that may preserve structure but are not optimized for identity clustering.
Furthermore, the proposed diversity regularization promotes decorrelation among embedding dimensions, resulting in a compact, well-conditioned representation with an effective rank matched to the number of discriminative factors. This combination of discriminative training and structured regularization produces clusters that are more compact and better separated, which is reflected across all evaluation metrics in Table 2.
@@
Figure 2. LFW & Olivetti Faces: t-SNE visualization of face embeddings. Left to right: Olivetti (original data), Olivetti (metric learning), LFW (original data), LFW (metric learning). Metric learning produces distinct, identity-based clusters.
The practical impact of this clustering is evident in nearest-neighbor retrieval. In Figures 3 and 4, the nearest neighbors of a query image in the metric learning space are consistently of the same person, despite variations in pose and lighting. The small distances between anchor and positive pairs confirm the model successfully compacts same-identity samples.
Figure 3. LFW Faces: Nearest-neighbor retrieval on the LFW dataset using the metric learning embedding. The model correctly identifies same-identity images as the closest neighbors, with small corresponding Euclidean distances.
Figure 4. Olivetti Faces: Nearest-neighbor retrieval on the Olivetti Faces dataset using the metric learning embedding. The results demonstrate robust, identity-based retrieval.
The rank sensitivity analysis reveals a fundamental limitation of fixed-rank tensor decompositions for semantic clustering tasks. On the challenging LFW dataset, CP, Tucker and t-SVD decompositions fail to achieve meaningful clustering across all ranks, with Silhouette scores remaining near zero or negative. This indicates that the low-rank constraints destroy the semantic structure necessary for identity separation. While Tucker and t-SVD decompositions show a slight improvement with higher ranks (from -0.0689 at R5 to -0.0037 at R20 for Tucker and from -0.0690 at R5 to -0.0037 at R20 for t-SVD), the gains are minimal. The Olivetti dataset, with its controlled conditions, shows more rank sensitivity, particularly for Tucker and t-SVD decompositions where performances improve substantially with higher ranks (Tucker: 0.1303 at R10 to 0.1845 at R20 similat to t-SVD). However, even the best fixed-rank result (Tucker-R20/t-SVD: 0.1845) is outperformed by the metric learning approach (0.8566), demonstrating that rank constraints inherently limit the ability to capture semantically meaningful representations, regardless of rank selection.